OpenStudy (anonymous):

What is the distance between points a(6,-1) and b(-2,-1)

OpenStudy (anonymous):

\[\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (anonymous):

|AB|=(-2,-1)-(6,-1)=(-8,0) 

OpenStudy (anonymous):

Then what iranmeah91 according to brother pathagoris thats nothing more than SQRT(-8)^2+(0)^2=8